OpenStudy (anonymous):

a dragster accelerates to a speed of 112 m/s over a distance of 398 m. determine the acceleration of the dragster. assume uniform acceleration.

OpenStudy (anonymous):

we can use eqn of motion...v^2=u^2+2*a*S...assume inital velocity(u) to be zero..then..v=112..and S=398...so u get a=15.759m/s^2

OpenStudy (anonymous):

Formula to be used are: \[a = (V _{f}^{2} +V _{i}^{1} ) \div 2dx\] \[t= dx/V _{1} + V _{2}\] Then the given values are: Initial Velocity= 0 Final Velocity= 112 m/s Distance= 398 Time=? 3.55 secs First, we get the time. Time= 398/ 0+ 112 Time= 3.55 secs Then we substitute. a= [(112 m/s^2)- (0^2)] / 2(398 m) a= 15.75 m/s^2 ** note that when squaring a value, the unit of the value must be squared too :)